# The cone conjecture for primitive symplectic varieties over a field of characteristic zero and an application

### v1

@@ 2.2. The Picard scheme and the Picard lattice of a primitive symplectic variety
In this subsection, we fix a field $F$ of characteristic $0$ and a projective primitive symplectic variety $X$ over $F$. We quickly recall the definition and some properties of the Picard scheme of $X$. This scheme will be useful for understanding the interaction between the Picard group of $X$ and the group of invariant elements for the action of the absolute Galois group $\mathop{\rm Gal}\nolimits(\overline{F}/F)$ on $\mathop{\rm Pic}\nolimits(X_{\overline{F}})$. We will use these results freely in the following without necessarily referring to them explicitly. For a more in-depth reading on the topic, see for example [undefw, Part. 5].
@@
Theorem 2.9 (Theorem 9.4.8 - [undefw]).
@@
Proposition 2.10 (Proposition 4.3.2 - [undefp]).
@@
Proposition 2.11 (Corollary 4.1.3 - [undefp]).
@@
Therefore, both $\mathop{\rm Pic}\nolimits(X)$ and $\mathop{\rm Pic}\nolimits(X_{\overline{F}})^{G_{F}}$ are finitely generated, and since the Brauer group $\mathop{\rm Br}\nolimits(F)$ is a torsion group ([undefp, Corollary 1.3.6]), $\mathop{\rm Pic}\nolimits(X)$ is of finite index in $\mathop{\rm Pic}\nolimits(X_{\overline{F}})^{G_{F}}$. As we will see in Proposition 2.16, we can define on $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}:=\mathop{\rm Pic}\nolimits(X)/\operatorname{(torsion)}$ a structure of lattice. We briefly note that there is a natural inclusion $\mathop{\rm Pic}\nolimits(X)\hookrightarrow\mathop{\rm\textbf{Pic}}\nolimits_{X/F}(F)$, which is in general not an equality; the failure of equality is measured by the group $\mathop{\rm Br}\nolimits(F)$ [undefp, Corollary 2.5.9].
Lemma 2.3 and Corollary 2.6 are particularly useful for defining a symmetric bilinear pairing on a second étale cohomology group of a symplectic primitive variety. Indeed, when working with the field $F=\mathbb{C}$ and $Y$ a primitive symplectic variety over $\mathbb{C}$, we can look at the associated analytic primitive symplectic variety $Y^{\operatorname{an}}$. For such varieties, we know that there exists an integral nondegenerate quadratic form on the second cohomology group similar to the Beauville-Bogomolov-Fujiki form (abbreviated as the BBF form in the following) on irreducible holomorphic symplectic manifolds (also known as hyperkähler manifolds). The existence of such a form is the result of the work of several mathematicians, see for example [undefd, $\S$5.1] for more information. Before giving the definition, we need a preliminary result.
@@
In particular, every reflexive two-form on $Y$ defines a unique class in $H^{2}(Y,\mathbb{C})$. With a slight abuse of notation, if $(Y,\sigma)$ is an analytic primitive symplectic variety, we will also denote by $\sigma$ the resulting cohomology class in $H^{2}(Y,\mathbb{C})$. Below, we provide the definition of the BBF form as defined by T. Kirschner in [undefan, Definition 3.2.7], but we note that there is also another definition given by Y. Namikawa in [undefaah, Theorem 8, (2)]. Fortunately, these two definitions are equivalent, see for example [undefaaq, Corollary 24] for a proof.
@@
Let $Y$ be an analytic primitive symplectic variety of dimension $2m$ and let $\sigma\in H^{2,0}(Y)$ be the class of a holomorphic symplectic $2$-form on $Y^{\mathop{\rm reg}\nolimits}$ satisfying $\int_{Y}(\sigma\cdot\overline{\sigma})^{m}=1$. Then the quadratic form $q_{Y}:H^{2}(Y,\mathbb{C})\to\mathbb{C}$ defined by
@@
does not depend on the choice of $\sigma$ and is called the Beauville-Bogomolov-Fujiki form (BBF form) on $Y$. It is nondegenerate and, up to scaling by a real number, is defined over $\mathbb{Z}$. Furthermore, the associated bilinear form has signature $(3,b_{2}(Y)-3)$, satisfies $q_{Y}|_{\left(H^{2,0}(Y)\oplus H^{0,2}(Y)\right)_{\mathbb{R}}}>0$ and the orthogonal complement to $H^{2,0}(Y)\oplus H^{0,2}(Y)$ (with respect to $q_{Y}$) equals $H^{1,1}(Y)$.
From now on, we will define an analogue of the BBF form on a projective primitive symplectic variety $X$ over a field $F$ of characteristic $0$.
@@
always splits. After choosing a section $s:\mathop{\rm Pic}\nolimits(X_{\overline{F}})_{\mathop{\rm tf}\nolimits}\hookrightarrow\mathop{\rm Pic}\nolimits(X_{\overline{F}})$, simply precompose $c_{1}^{\ell}:\mathop{\rm Pic}\nolimits(X_{\overline{F}})\otimes_{\mathbb{Z}}\mathbb{Z}_{\ell}\hookrightarrow H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X_{\overline{F}},\mathbb{Z}_{\ell})$ with $s\otimes 1_{\mathbb{Z}_{\ell}}$ to obtain an injection. We preferred to base our proof without the choice of such a section, even if this meant that our statement did not cover all the prime integers.
@@
Let $F$ be a field of characteristic $0$, $G_{F}:=\mathop{\rm Gal}\nolimits(\overline{F}/F)$ the absolute Galois group of $F$ and $X$ a primitive symplectic variety over $F$ of dimension $2m$. Then, for all but finitely many prime number $\ell\geq 2$, there exists a unique $G_{F}$-invariant pairing $q_{X,\ell}:H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X_{\overline{F}},\mathbb{Z}_{\ell})^{\otimes 2}\to\mathbb{Z}_{\ell}$ such that
@@
    For any $\alpha_{1},\dotsb,\alpha_{2m}\in H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X_{\overline{F}},\mathbb{Z}_{\ell})$:
@@
By Corollary 2.6, we choose a finitely generated subfield $F^{\prime}\subset F$ over $\mathbb{Q}$ and $X^{\prime}$ a primitive symplectic variety over $F^{\prime}$ such that $X=X^{\prime}_{F}$. By fixing an embedding $\overline{F^{\prime}}\subset\mathbb{C}$, we know from Lemma 2.3 that $X^{\prime}_{\mathbb{C}}$ is a primitive symplectic variety over $\mathbb{C}$. By [undefaac, VI - Corollary 4.3], for any prime integer $\ell$, we have $H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X_{\overline{F}},\mathbb{Z}_{\ell})\simeq H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X^{\prime}_{\overline{F^{\prime}}},\mathbb{Z}_{\ell})\simeq H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X^{\prime}_{\mathbb{C}},\mathbb{Z}_{\ell})$ and by the comparison theorem [undefb, Theorem 2], we obtain an isomorphism $H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X^{\prime}_{\overline{F^{\prime}}},\mathbb{Z}_{\ell})\simeq H^{2}(X^{\prime}_{\mathbb{C}}(\mathbb{C}),\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{Z}_{\ell}$. We define $q_{X,\ell}$ thanks to this isomorphism and Proposition 2.13. The rest of the proof is then identical to that of [undefaay, Proposition 2.1.5] where the key argument is that the BBF form on an irreducible holomorphic symplectic manifold satisfies Fujiki relation, a relation which we also have for complex primitive symplectic varieties, see [undefaaq, Theorem 2]. In particular, the definition of $q_{X,\ell}$ is independent of the choice of an embedding $\overline{F^{\prime}}\subset\mathbb{C}$ and this form verifies the properties of the statement. ∎
@@ 3. Absolute cone conjectures for terminal Q-factorial primitive symplectic varieties
In this section, we prove the nef and movable cone conjectures for primitive symplectic varieties over a field $F$ of characteristic $0$ with $\mathbb{Q}$-factorial and terminal singularities and with a second Betti number greater than or equal to $5$. Since these varieties are integral by definition, we will slightly abuse notation by identifying the notion of classes of Cartier divisors (for the linear equivalence relation) with the notion of classes of line bundles (for the isomorphism relation).
@@ 3.1. Preliminaries
Let $X$ be a primitive symplectic variety over a field $F$ of characteristic $0$ and let $\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}:=\mathop{\rm Pic}\nolimits(X)\otimes_{\mathbb{Z}}\mathbb{R}$ be the real Picard group of $X$. Inside this vector space, we consider the following convex cones.
@@
Alternatively, we may work with the Neron-Severi space instead of the Picard space. Indeed, if $X$ is a primitive symplectic variety over a field $F$ of characteristic $0$, then $\mathop{\rm Pic}\nolimits(X)_{\mathbb{K}}\simeq N^{1}(X)_{\mathbb{K}}$ for $\mathbb{K}=\mathbb{Q}$ or $\mathbb{R}$ by [undefw, Corollary 9.5.13].
@@
Finally, we show that $D$ is movable if and only if $D_{L}$ is movable. We denote by $\mathop{\rm Bs}\nolimits(D)$ and $\mathop{\rm Bs}\nolimits(D_{L})$ the base loci of $\left\lvert D\right\rvert$ and $\left\lvert D_{L}\right\rvert$ respectively. Since we are only interested in the codimension, we equip these schemes with their reduced structure. We have the set-theoretic equality $p^{-1}\left(\mathop{\rm Bs}\nolimits(D)\right)=\mathop{\rm Bs}\nolimits(D_{L})$ since $H^{0}(X_{L},D_{L})=H^{0}(X,D)\otimes_{F}L$. Consider the base change $q:\mathop{\rm Bs}\nolimits(D)\otimes_{F}L\to\mathop{\rm Bs}\nolimits(D)$. The map $Z\mapsto\overline{q(Z)}$ is a well-defined surjective map from the set of irreducible components of $\mathop{\rm Bs}\nolimits(D)\otimes_{F}L$ to the set of irreducible components of $\mathop{\rm Bs}\nolimits(D)$, and $\dim Z=\dim\overline{p(Z)}$, see for example [undefaa, Corollary 5.45, Exercise 5.12]. Since $X$ and $X_{L}$ are irreducible, we have
@@
by [undefaa, Proposition 5.30.(2)]. Since $\dim X=\dim X_{L}$, it follows that $\operatorname{codim}_{X_{L}}Z=\operatorname{codim}_{X}\overline{p(Z)}$ for every irreducible components $Z$ of $\mathop{\rm Bs}\nolimits(D)\otimes_{F}L$. As the underlying topological space of $\mathop{\rm Bs}\nolimits(D)\otimes_{F}L$ is $p^{-1}\left(\mathop{\rm Bs}\nolimits(D)\right)=\mathop{\rm Bs}\nolimits(D_{L})$, we deduce that
@@
We only prove the case of ampleness; the other cases are treated in a similar manner. Let $x=\sum a_{i}x_{i}$ be an element of $\mathop{\rm Amp}\nolimits(\overline{X})\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$, where the $a_{i}$ are real numbers and the $x_{i}$ are integral classes of ample line bundles on $\overline{X}$. Let us take a finite Galois extension $F\subset F^{\prime}$ such that the $x_{i}$ are contained in $\mathop{\rm Pic}\nolimits(X_{F^{\prime}})_{\mathop{\rm tf}\nolimits}$ and let $\Gamma:=\mathop{\rm Gal}\nolimits(F^{\prime}/F)$. For each index $i$, we consider
@@
Lemma 3.5 (Lemma 3.8 - [undefm]).
@@
Lemma 3.5 is stated in [undefm] for closed convex cones, but it remains true more generally for convex cones $C$ satisfying the assumption $S\cap\operatorname{int}(C)\neq\emptyset$. Also, note that the proof of Lemma 3.4 shows more generally that
@@
Let $X$ be a primitive symplectic variety over an algebraically closed field $L$ of characteristic $0$. A prime Weil divisor $E$ on $X$ is called a prime exceptional divisor if it is $\mathbb{Q}$-Cartier and satisfies $q_{X}(E,E)<0$. For any such prime exceptional divisor $E$, we can associate the reflection
@@
Lemma 3.8 (Lemma 2.15 - [undefx]).
@@
The induced isomorphism $p^{*}_{\mathop{\rm tf}\nolimits}:\mathop{\rm Pic}\nolimits(X_{M})_{\mathop{\rm tf}\nolimits}\xrightarrow{\simeq}\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ is an isometry with respect to the BBF forms on $X$ and $X_{M}$.
@@
By [undefw, Corollary 9.5.13] and Lemma 2.7, the Picard scheme $\mathop{\rm\textbf{Pic}}\nolimits_{X/L}$ is a smooth group scheme over $L$ of dimension $h^{0}(X,{\mathcal{T}}_{X})=0$, and is therefore étale over $L$. Since $L$ is algebraically closed, we deduce that $\mathop{\rm\textbf{Pic}}\nolimits_{X/L}$ is a disjoint union of copies of $\mathop{\rm Spec}\nolimits L$, i.e. is a constant group scheme. In particular, the inclusion $\iota:L\hookrightarrow M$ induces an isomorphism
$\mathop{\rm\textbf{Pic}}\nolimits_{X/L}(\iota):\mathop{\rm\textbf{Pic}}\nolimits_{X/L}(\mathop{\rm Spec}\nolimits L)\xrightarrow{\simeq}\mathop{\rm\textbf{Pic}}\nolimits_{X/L}(\mathop{\rm Spec}\nolimits M).$ (3.2)
Similarly, using the fact that $\mathop{\rm\textbf{Pic}}\nolimits_{X_{M}/M}=\mathop{\rm\textbf{Pic}}\nolimits_{X/L}\otimes_{L}M$, it is clear that this is also a constant group scheme such that $\mathop{\rm\textbf{Pic}}\nolimits_{X_{M}/M}(M)=\mathop{\rm\textbf{Pic}}\nolimits_{X/L}(M)$. By [undefp, Corollary 2.5.8], we have $\mathop{\rm\textbf{Pic}}\nolimits_{X/L}(L)\simeq\mathop{\rm Pic}\nolimits(X)$ and $\mathop{\rm\textbf{Pic}}\nolimits_{X_{M}/M}(M)\simeq\mathop{\rm Pic}\nolimits(X_{M})$. We then observe that the isomorphism (3.2) then corresponds to
$p^{*}:\mathop{\rm Pic}\nolimits(X)\xrightarrow{\simeq}\mathop{\rm Pic}\nolimits(X_{M}).$
The fact that $p^{*}_{\mathop{\rm tf}\nolimits}$ is an isometry is a consequence both of the description of this isomorphism as the pullback of line bundles, and of the construction of the BBF form through natural and functorial isomorphisms, and this concludes the proof. ∎

### v2

@@ 2.2. The Picard scheme and the Picard lattice of a primitive symplectic variety
In this subsection, we fix a field $F$ of characteristic $0$ and a projective primitive symplectic variety $X$ over $F$. We quickly recall the definition and some properties of the Picard scheme of $X$. This scheme will be useful for understanding the interaction between the Picard group of $X$ and the group of invariant elements under the action of the absolute Galois group $\mathop{\rm Gal}\nolimits(\overline{F}/F)$ on $\mathop{\rm Pic}\nolimits(X_{\overline{F}})$. We will use these results freely in the following without necessarily referring to them explicitly. For a more in-depth reading on the topic, see for example [undefv, Part. 5].
@@
Theorem 2.9 (Theorem 9.4.8 - [undefv]).
@@
Proposition 2.10 (Proposition 4.3.2 - [undefo]).
@@
Proposition 2.11 (Corollary 4.1.3 - [undefo]).
@@
Therefore, both $\mathop{\rm Pic}\nolimits(X)$ and $\mathop{\rm Pic}\nolimits(X_{\overline{F}})^{G_{F}}$ are finitely generated, and since the Brauer group $\mathop{\rm Br}\nolimits(F)$ is a torsion group ([undefo, Corollary 1.3.6]), $\mathop{\rm Pic}\nolimits(X)$ is of finite index in $\mathop{\rm Pic}\nolimits(X_{\overline{F}})^{G_{F}}$. As we will see in Proposition 2.16, we can define on $\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}:=\mathop{\rm Pic}\nolimits(X)/\operatorname{(torsion)}$ a structure of lattice. We briefly note that there is a natural inclusion $\mathop{\rm Pic}\nolimits(X)\hookrightarrow\mathop{\rm\textbf{Pic}}\nolimits_{X/F}(F)$, which is in general not an equality; the failure of equality is measured by the group $\mathop{\rm Br}\nolimits(F)$ [undefo, Corollary 2.5.9].
Lemma 2.3 and Corollary 2.6 are particularly useful for defining a symmetric bilinear pairing on a second étale cohomology group of a primitive symplectic variety. Indeed, when working with the field $F=\mathbb{C}$ and $Y$ a primitive symplectic variety over $\mathbb{C}$, we can look at the associated analytic primitive symplectic variety $Y^{\operatorname{an}}$. For such varieties, we know that there exists an integral nondegenerate quadratic form on the second cohomology group similar to the Beauville-Bogomolov-Fujiki form on irreducible holomorphic symplectic manifolds (also known as hyperkähler manifolds). The existence of such a form is the result of the work of several mathematicians, see for example [undefd, $\S$5.1] for more information. Before giving the definition, we need a preliminary result.
@@
In particular, every reflexive two-form on $Y$ defines a unique class in $H^{2}(Y,\mathbb{C})$. Since the singularities of a complex projective primitive symplectic variety $(Y,\sigma)$ are canonical by Lemma 2.7, we will also denote by $\sigma$ the resulting cohomology class in $H^{2}(Y,\mathbb{C})$ for simplicity. Below, we provide the definition of the Beauville-Bogomolov-Fujiki form as defined by T. Kirschner in [undefam, Definition 3.2.7], but we note that there is also another definition given by Y. Namikawa in [undefaah, Theorem 8, (2)]. Fortunately, these two definitions are equivalent, see for example [undefaaq, Corollary 24] for a proof.
@@
Let $Y$ be a complex projective primitive symplectic variety of dimension $2m$ and let $\sigma\in H^{2,0}(Y)$ be the class of a holomorphic symplectic $2$-form on $Y^{\mathop{\rm reg}\nolimits}$ satisfying $\int_{Y}(\sigma\cdot\overline{\sigma})^{m}=1$. Then the quadratic form $q_{Y}:H^{2}(Y,\mathbb{C})\to\mathbb{C}$ defined by
@@
does not depend on the choice of $\sigma$ and is called the Beauville-Bogomolov-Fujiki form on $Y$. It is nondegenerate and, up to scaling by a real number, is defined over $\mathbb{Z}$. Furthermore, the associated bilinear form has signature $(3,b_{2}(Y)-3)$, satisfies $q_{Y}|_{\left(H^{2,0}(Y)\oplus H^{0,2}(Y)\right)_{\mathbb{R}}}>0$ and the orthogonal complement to $H^{2,0}(Y)\oplus H^{0,2}(Y)$ with respect to $q_{Y}$ is equal to $H^{1,1}(Y)$.
From now on, we will define an analogue of the Beauville-Bogomolov-Fujiki form on a projective primitive symplectic variety $X$ over a field $F$ of characteristic $0$.
@@
always splits. After choosing a section $s$, simply precompose $c_{1}^{\ell}:\mathop{\rm Pic}\nolimits(X_{\overline{F}})\otimes_{\mathbb{Z}}\mathbb{Z}_{\ell}\hookrightarrow H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X_{\overline{F}},\mathbb{Z}_{\ell})$ with $s\otimes 1$ to obtain an injection. We preferred to base our proof without the choice of such a section, even if this meant that our statement did not cover all the prime integers.
@@
Let $F$ be a field of characteristic $0$, $G_{F}:=\mathop{\rm Gal}\nolimits(\overline{F}/F)$ the absolute Galois group of $F$ and $X$ a primitive symplectic variety over $F$ of dimension $2m$. Then, for all but finitely many prime number $\ell\geq 2$, there exists a unique $G_{F}$-invariant pairing $q_{X,\ell}:H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X_{\overline{F}},\mathbb{Z}_{\ell})^{\otimes 2}\to\mathbb{Z}_{\ell}$ satisfying the following properties.
@@
    For any $\alpha_{1},\dotsb,\alpha_{2m}\in H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X_{\overline{F}},\mathbb{Z}_{\ell})$, we have
@@
By Corollary 2.6, choose a finitely generated subfield $F^{\prime}\subset F$ over $\mathbb{Q}$ and $X^{\prime}$ a primitive symplectic variety over $F^{\prime}$ such that $X=X^{\prime}_{F}$. By fixing an embedding $\overline{F^{\prime}}\subset\mathbb{C}$, we know from Lemma 2.3 that $X^{\prime}_{\mathbb{C}}$ is a primitive symplectic variety over $\mathbb{C}$. By [undefaac, VI - Corollary 4.3], for any prime integer $\ell$, we have $H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X_{\overline{F}},\mathbb{Z}_{\ell})\simeq H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X^{\prime}_{\overline{F^{\prime}}},\mathbb{Z}_{\ell})\simeq H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X^{\prime}_{\mathbb{C}},\mathbb{Z}_{\ell})$ and by the comparison theorem [undefb, Theorem 2], we obtain an isomorphism $H^{2}_{\mathop{\rm\acute{e}t}\nolimits}(X^{\prime}_{\overline{F^{\prime}}},\mathbb{Z}_{\ell})\simeq H^{2}(X^{\prime\operatorname{an}}_{\mathbb{C}},\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{Z}_{\ell}$. We define $q_{X,\ell}$ thanks to this isomorphism and Proposition 2.13. The rest of the proof is then identical to that of [undefaay, Proposition 2.1.5] where the key argument is that the Beauville-Bogomolov-Fujiki form on an irreducible holomorphic symplectic manifold satisfies Fujiki relation, a relation which we also have for complex primitive symplectic varieties, see [undefaaq, Theorem 2]. In particular, the definition of $q_{X,\ell}$ is independent of the choice of an embedding $\overline{F^{\prime}}\subset\mathbb{C}$ and this form verifies the properties of the statement. ∎
@@ 3. Absolute cone conjectures for terminal Q-factorial primitive symplectic varieties
In this section, we prove the movable and nef cone conjectures for a projective primitive symplectic variety $X$ over a field $F$ of characteristic $0$ with $b_{2}(X)\geq 5$ and $\mathbb{Q}$-factorial terminal singularities. Since $X$ is integral by definition, we will slightly abuse notation by identifying the notion of classes of Cartier divisors (for the linear equivalence relation) with the notion of classes of line bundles (for the isomorphism relation).
@@ 3.1. Preliminaries
Let $X$ be a primitive symplectic variety over a field $F$ of characteristic $0$ and let $\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}:=\mathop{\rm Pic}\nolimits(X)\otimes_{\mathbb{Z}}\mathbb{R}$ be the real Picard space of $X$. Inside this vector space, we consider the following convex cones.
@@
Alternatively, we may work with the Neron-Severi space instead of the Picard space. Indeed, if $X$ is a primitive symplectic variety over a field $F$ of characteristic $0$, then $\mathop{\rm Pic}\nolimits(X)_{\mathbb{K}}\simeq N^{1}(X)_{\mathbb{K}}$ for $\mathbb{K}=\mathbb{Q}$ or $\mathbb{R}$ by [undefv, Corollary 9.5.13].
@@
Finally, we show that $D$ is movable if and only if $D_{L}$ is movable. We denote by $\mathop{\rm Bs}\nolimits(D)$ and $\mathop{\rm Bs}\nolimits(D_{L})$ the base loci of $\left\lvert D\right\rvert$ and $\left\lvert D_{L}\right\rvert$ respectively. Since we are only interested in the codimension, we equip these schemes with their reduced structure. We have the set-theoretic equality $p^{-1}\left(\mathop{\rm Bs}\nolimits(D)\right)=\mathop{\rm Bs}\nolimits(D_{L})$ since $H^{0}(X_{L},D_{L})=H^{0}(X,D)\otimes_{F}L$. Consider the base change $q:\mathop{\rm Bs}\nolimits(D)\otimes_{F}L\to\mathop{\rm Bs}\nolimits(D)$. The map $Z\mapsto\overline{q(Z)}$ is a well-defined surjective map from the set of irreducible components of $\mathop{\rm Bs}\nolimits(D)\otimes_{F}L$ to the set of irreducible components of $\mathop{\rm Bs}\nolimits(D)$, and $\dim Z=\dim\overline{p(Z)}$, see for example [undefz, Corollary 5.45, Exercise 5.12]. Since $X$ and $X_{L}$ are irreducible, we have
@@
by [undefz, Proposition 5.30.(2)]. Since $\dim X=\dim X_{L}$, it follows that $\operatorname{codim}_{X_{L}}Z=\operatorname{codim}_{X}\overline{p(Z)}$ for every irreducible components $Z$ of $\mathop{\rm Bs}\nolimits(D)\otimes_{F}L$. As the underlying topological space of $\mathop{\rm Bs}\nolimits(D)\otimes_{F}L$ is $p^{-1}\left(\mathop{\rm Bs}\nolimits(D)\right)=\mathop{\rm Bs}\nolimits(D_{L})$, we deduce that
@@
We only prove the case of ampleness, the other cases are treated in a similar manner. Let $x=\sum a_{i}x_{i}$ be an element of $\mathop{\rm Amp}\nolimits(\overline{X})\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$, where the $a_{i}$ are real numbers and the $x_{i}$ are integral classes of ample line bundles on $\overline{X}$. Take a finite Galois extension $F\subset F^{\prime}$ such that the $x_{i}$ are contained in $\mathop{\rm Pic}\nolimits(X_{F^{\prime}})_{\mathop{\rm tf}\nolimits}$ and let $\Gamma:=\mathop{\rm Gal}\nolimits(F^{\prime}/F)$. For each index $i$, we consider
@@
Lemma 3.5 (Lemma 3.8 - [undefl]).
@@
Lemma 3.5 is stated in [undefl] for closed convex cones, but it remains true more generally for convex cones $C$ satisfying the assumption $S\cap\operatorname{int}(C)\neq\emptyset$. Also, note that the proof of Lemma 3.4 shows more generally that
@@
Let $X$ be a primitive symplectic variety over an algebraically closed field $L$ of characteristic $0$. A prime Weil divisor $E$ on $X$ is called a prime exceptional divisor if it is $\mathbb{Q}$-Cartier and satisfies $q_{X}(E,E)<0$. For any such prime exceptional divisor $E$, we define the associated reflection
@@
Lemma 3.8 (Lemma 2.15 - [undefw]).
@@
The induced isomorphism $p^{*}_{\mathop{\rm tf}\nolimits}:\mathop{\rm Pic}\nolimits(X_{M})_{\mathop{\rm tf}\nolimits}\xrightarrow{\simeq}\mathop{\rm Pic}\nolimits(X)_{\mathop{\rm tf}\nolimits}$ is an isometry with respect to the Beauville-Bogomolov-Fujiki forms on $X$ and $X_{M}$.
@@
By [undefv, Corollary 9.5.13] and Lemma 2.7, the Picard scheme $\mathop{\rm\textbf{Pic}}\nolimits_{X/L}$ is a smooth group scheme over $L$ of dimension $h^{0}(X,{\mathcal{T}}_{X})=0$, and is therefore étale over $L$. Since $L$ is algebraically closed, we deduce that $\mathop{\rm\textbf{Pic}}\nolimits_{X/L}$ is a disjoint union of copies of $\mathop{\rm Spec}\nolimits L$, i.e. is a constant group scheme. In particular, the inclusion $\iota:L\hookrightarrow M$ induces an isomorphism
$\mathop{\rm\textbf{Pic}}\nolimits_{X/L}(\iota):\mathop{\rm\textbf{Pic}}\nolimits_{X/L}(L)\xrightarrow{\simeq}\mathop{\rm\textbf{Pic}}\nolimits_{X/L}(M).$ (3.2)
Similarly, using the fact that $\mathop{\rm\textbf{Pic}}\nolimits_{X_{M}/M}=\mathop{\rm\textbf{Pic}}\nolimits_{X/L}\otimes_{L}M$, it is clear that $\mathop{\rm\textbf{Pic}}\nolimits_{X_{M}/M}$ is also a constant group scheme such that $\mathop{\rm\textbf{Pic}}\nolimits_{X_{M}/M}(M)=\mathop{\rm\textbf{Pic}}\nolimits_{X/L}(M)$. By [undefo, Corollary 2.5.8], we have $\mathop{\rm\textbf{Pic}}\nolimits_{X/L}(L)\simeq\mathop{\rm Pic}\nolimits(X)$ and $\mathop{\rm\textbf{Pic}}\nolimits_{X_{M}/M}(M)\simeq\mathop{\rm Pic}\nolimits(X_{M})$. We then observe that the isomorphism (3.2) corresponds to the pullback morphism $p^{*}:\mathop{\rm Pic}\nolimits(X)\to\mathop{\rm Pic}\nolimits(X_{M})$. The fact that $p^{*}_{\mathop{\rm tf}\nolimits}$ is an isometry is a consequence both of the description of this isomorphism as the pullback of line bundles, and of the construction of the Beauville-Bogomolov-Fujiki form through natural and functorial isomorphisms, and this concludes the proof. ∎